# Benchmarking Retrieval-Augmented Generation for Medicine

### v1

@@ 1 Introduction
Large Language Models (LLMs) have revolutionized the way people seek information online, from searching to directly asking chatbots for answers. Although recent studies have shown their state-of-the-art capabilities of question answering (QA) in both general and medical domains (OpenAI et al., 2023; Anil et al., 2023; Touvron et al., 2023b; Singhal et al., 2023a; Nori et al., 2023a), LLMs often generate plausible-sounding but factually incorrect responses, commonly known as hallucination Ji et al. (2023). Additionally, the training corpora of LLMs might not include the latest knowledge, such as recent updates of clinical guidelines. These issues can be especially dangerous in high-stakes domains such as healthcare Tian et al. (2024).
@@
Based on the Mirage benchmark, we systematically evaluated different MedRag solutions and studied the effects of each RAG component on the final performance from a multidimensional perspective. For various LLMs, there is a 1% to 18% relative performance increase by using MedRag compared to chain-of-thought prompting (Wei et al., 2022). Notably, with MedRag, GPT-3.5 and Mixtral Jiang et al. (2024) can achieve comparable performance to GPT-4 OpenAI et al. (2023) on Mirage. On the corpus dimension, we found different tasks have a preference over the retrieval corpus. While point-of-care articles and textbooks are solely helpful for examination questions, PubMed is a robust choice for all Mirage tasks. Our experiments also demonstrate that a combination of all corpora can be a more comprehensive choice. On the retriever dimension, BM25 Robertson et al. (2009) and the domain-specific MedCPT Jin et al. (2023a) retriever display superior performance on our Mirage benchmark. The performance can be further enhanced by combining results from multiple retrievers. Beyond the evaluation results on Mirage, we found a log-linear scaling relationship between model performance and the number of retrieved snippets. We also observed a “lost-in-the-middle” phenomenon (Liu et al., 2023) between model performance and the position of the ground-truth snippet. Finally, we provide several practical recommendations based on the results and analyses, which can guide the application and future research of RAG in the biomedical domain.
@@ 2.1 Retrieval-augmented Generation
In biomedicine, there have also been various explorations on how LLMs can improve literature information-seeking and clinical decision-making with RAG Frisoni et al. (2022); Naik et al. (2022); Jin et al. (2023b); Lála et al. (2023); Zakka et al. (2024); Jeong et al. (2024), but their evaluations are not comprehensive. Nevertheless, current systematic evaluations in biomedicine typically focus on the vanilla LLMs without RAG (Chen et al., 2023a; Nori et al., 2023a). Our study provides the first systematic evaluations of RAG systems in medicine.
@@ 7 Conclusion
While our study provides systematic evaluations and practical recommendations for medical RAG systems, there are several limitations that need to be acknowledged. First, there have been novel developments in the architecture of RAG (e.g., active RAG, Jiang et al., 2023). However, we mainly evaluate the vanilla RAG architecture where the retrieved documents are directly prepended in the LLM context because this is the most widely implemented architecture. Evaluating new RAG system designs remains an important direction to explore. Second, while the coverage of corpora, retrievers, and LLMs in MedRag is reasonably comprehensive, there are other potentially useful resources that can also be incorporated into MedRag in future work, such as the full-text articles from PubMed Central (PMC)555https://www.ncbi.nlm.nih.gov/pmc/ and Frequently Asked Questions (FAQs) from trustworthy sources (Ben Abacha and Demner-Fushman, 2019). Third, we only evaluate the retrieval component for PubMedQA* and BioASQ-Y/N since the other three examination datasets lack labels of ground-truth supporting documents. Further research should also evaluate whether the retrieved snippets are actually helpful for the examination datasets, and explore the use of cross-encoder re-rankers to improve the retrieval performance for relevant information. Fourth, while QA is the most commonly used task for evaluating biomedical LLMs, there are also other knowledge-intensive tasks that might benefit from MedRag, such as claim verification (Wadden et al., 2020). Following most other studies, we use the format of multi-choice questions for large-scale and automatic evaluation of medical QA. Although we restrict the retrieval phase to having no access to the choices, LLMs still need to use them as input for the final prediction. The rationales generated by MedRag remain to be evaluated as well. As the goal of this study is to systematically benchmark the most commonly used medical RAG settings, we leave the potential solutions of the above-mentioned limitations to future work.
@@ B.1 Document Collections
As a large-scale open-source encyclopedia, Wikipedia is frequently used as a corpus in information retrieval tasks Thakur et al. (2021). We select Wikipedia as one of the corpora to see if the general domain database can be used to improve the ability of medical QA. We download the processed Wikipedia data from HuggingFace171717https://huggingface.co/datasets/wikipedia and also chunked the text with LangChain.

### v2

@@ 1 Introduction
Large Language Models (LLMs) have revolutionized the way people seek information online, from searching to directly asking chatbots for answers. Although recent studies have shown their state-of-the-art capabilities of question answering (QA) in both general and medical domains (OpenAI et al., 2023; Anil et al., 2023; Touvron et al., 2023b; Singhal et al., 2023a; Nori et al., 2023a), LLMs often generate plausible-sounding but factually incorrect responses, commonly known as hallucination Ji et al. (2023). Also, the training corpora of LLMs might not include the latest knowledge, such as recent updates of clinical guidelines. These issues can be especially dangerous in high-stakes domains such as healthcare Tian et al. (2024); Hersh (2024).
@@
Based on the Mirage benchmark, we systematically evaluated different MedRag solutions and studied the effects of each component on overall performance from a multidimensional perspective. For various LLMs, there is a 1% to 18% relative performance increase using MedRag compared to chain-of-thought prompting (Wei et al., 2022). Notably, with MedRag, GPT-3.5 and Mixtral Jiang et al. (2024) can achieve comparable performance to GPT-4 OpenAI et al. (2023) on Mirage. On the corpus dimension, we found different tasks have a preference over the retrieval corpus. While point-of-care articles and textbooks are solely helpful for examination questions, PubMed is a robust choice for all Mirage tasks. Our results also show that a combination of all corpora can be a more comprehensive choice. On the retriever dimension, BM25 Robertson et al. (2009) and the domain-specific MedCPT Jin et al. (2023a) retriever display superior performance on our Mirage benchmark. The performance can be further enhanced by combining multiple retrievers. Beyond the evaluation results on Mirage, we found a log-linear scaling relationship between model performance and the number of retrieved snippets. We also observed a “lost-in-the-middle” phenomenon (Liu et al., 2023) between model performance and the position of the ground-truth snippet. Finally, we provide several practical recommendations based on the results and analyses, which can guide the application and future research of RAG in the biomedical domain.
@@ 2.1 Retrieval-augmented Generation
In biomedicine, there have also been various explorations on how LLMs can improve literature information-seeking and clinical decision-making with RAG Frisoni et al. (2022); Naik et al. (2022); Jin et al. (2023b); Lála et al. (2023); Zakka et al. (2024); Jeong et al. (2024); Wang et al. (2023), but their evaluations are not comprehensive. Nevertheless, current systematic evaluations in biomedicine typically focus on the vanilla LLMs without RAG (Chen et al., 2023a; Nori et al., 2023a). Our study provides the first systematic evaluations of RAG systems in medicine.
@@ 7 Conclusion
While our study provides systematic evaluations and practical recommendations for medical RAG systems, there are several limitations that need to be acknowledged. First, there have been novel developments in the architecture of RAG (e.g., active RAG, Jiang et al., 2023). However, we mainly evaluate the vanilla RAG architecture where the retrieved documents are directly prepended in the LLM context because this is the most widely implemented architecture. Evaluating new RAG system designs remains an important direction to explore. Second, while the coverage of corpora, retrievers, and LLMs in MedRag is reasonably comprehensive, there are other potentially useful resources that can also be incorporated into MedRag in future work, such as the full-text articles from PubMed Central (PMC)555https://www.ncbi.nlm.nih.gov/pmc/ and Frequently Asked Questions (FAQs) from trustworthy sources (Ben Abacha and Demner-Fushman, 2019). Third, we only evaluate the retrieval component for PubMedQA* and BioASQ-Y/N since the other three examination datasets lack labels of ground-truth supporting documents. Further research should also evaluate whether the retrieved snippets are actually helpful for the examination datasets, and explore the use of cross-encoder re-rankers to improve the retrieval performance for relevant information. Fourth, while QA is the most commonly used task for evaluating biomedical LLMs, there are also other knowledge-intensive tasks that might benefit from MedRag, such as claim verification (Wadden et al., 2020; Liu et al., 2024). Following most other studies, we use the format of multi-choice questions for large-scale and automatic evaluation of medical QA. Although we restrict the retrieval phase to having no access to the choices, LLMs still need to use them as input for the final prediction. The rationales generated by MedRag remain to be evaluated as well. As the goal of this study is to systematically benchmark the most commonly used medical RAG settings, we leave the potential solutions of the above-mentioned limitations to future work.
@@ B.1 Document Collections
As a large-scale open-source encyclopedia, Wikipedia is frequently used as a corpus in information retrieval tasks Thakur et al. (2021). We select Wikipedia as one of the corpora to see if the general domain database can be used to improve the ability of medical QA. We downloaded the processed Wikipedia data from HuggingFace171717https://huggingface.co/datasets/wikipedia and also chunked the text with LangChain.